OpenStudy (anonymous):

9,11,7 1,15,14 6,7,10 How do i classify the triangles as right acute or obtuse

OpenStudy (anonymous):

I'm assuming the numbers correspond to side lengths. In that case you could give them units and draw them out on a sheet of paper with a ruler (I would recommend centimeters if you are going to do this approach). Then, assuming that an angle is clearly either greater than 90(deg), less than 90(deg) or 90(deg) you could classify them that way. The other approach would be to order each triangles lengths from least to greatest. Then, knowing that the Pythagorean Theorem states that the square of the hypotenuse of a right triangle is equal to the sum of the squares of the "legs" of the triangle \[c^2=b^2+a^2\]you can determine whether the given side lengths correspond to an acute, right or obtuse triangle. If the summation of the squares of the legs is less than the square of the hypotenuse, the triangle is an acute triangle. If it is greater than, it is obtuse, and if it is equal to, it is a right triangle. From your problem, we can look at the first one as an example: \[9^2 + 7^2 = 130 > 121 = 11^2\] So the first triangle is an obtuse triangle. Again, to check this, you can assign a unit length such as centimeters and check the answer by drawing the triangle. Hope this helps!